OpenStudy (anonymous):

How to find a replacement value of resistance of several resistors (exactly 12 resistor that they have same value R) are arranged in a cube form?? thanks.

OpenStudy (anonymous):

It depends on the relative position of two points. R(adjacent)=11/19R R(diagonal of same face)=3/4R R(opposite corners)=5/6R Hint:symmetry